A refrigerator room has a wall 6m long and 3m high. The wall is built of 120mm thick brick, insulated on the inside with an 80mm layer of cork faced with a thin metal sheet. The coefficient of thermal conductivity of the brick is 1.15W/mK while that of the cork is 0.043W/mK. The exterior brick surface temperature of the wall is 21 degrees Celcius and that of the interior metal faced surface is -4d egrees Celcius throughout. Estimate the heat leakage through the wall in 24 hours and also the temperature of the interface between the cork and brick. 


1
Expert's answer
2021-11-29T00:17:12-0500

heat leakage

"Q = \frac{\triangle T}{R_{th}}=\frac{21+4}{1.15+0.043} = 20.95 W"